# Gil Ayache*, Menachem Pirchi*, Aviv Navon, Aviv Shamsian, Gill Hetz, Joseph Keshet aiOla Research, Israel

### v1

@@ III-A Zero-shot Open Type NER
We first evaluate WhisperNER on the three open-type NER benchmarks, namely VoxPopuli-NER, LibriSpeech-NER and Fleurs-NER. Each test example is annotated with positive NER tags based on the annotation scheme of [9]. To construct negative samples, we randomly sample negative NER labels from the entire dataset. We ensure that the number of negative NER labels matches the number of positive labels, resulting in a balanced dataset. The results are presented in Table I. both WhisperNER variants outperform all baseline methods in terms of NER F1 score, with only a slight decrease in WER, while being more parameter-efficient.
III-B Zero-shot NER
Next, we evaluate the zero-shot NER performance of our WhisperNER and WhisperNER-BIO models on the closed set entity sets of the MIT NER benchmarks [18]. The results are summarized in Table II. In this setup, WhisperNER does not perform as well as the baselines in terms of NER F1, showing a higher tendency to incorrectly tag non-entity spans. Recognizing these limitations, we identify areas for refinement and believe that a more in-depth analysis, and further enhancements, hold potential for significant improvements. We leave these pursuits for future research, focusing on optimizing WhisperNER to better handle closed set zero-shot scenarios.
III-C Supervised finetuning
Finally, we evaluate the supervised fine-tuning performance of our WhisperNER model on common NER benchmarks, with results summarized in Table III. For this experiment, the methods were fine-tuned on labeled datasets to assess their ability to recognize and tag entities in a supervised setting. For the baseline methods, we independently finetune both the Whipser model, and the NER baseline. Our method outperforms all baselines, achieving the best performance in both WER and NER F1 scores. These improvements demonstrate the effectiveness of our approach in jointly handling transcription and entity recognition tasks, further validating WhisperNER’s capability to generalize to structured NER datasets in a supervised setting.
@@ III-C Supervised finetuning
TABLE IV: Effect of negative sampling approach: We evaluate WhisperNER models trained using the NuNER dataset for 25K steps, on the zero-shot task over the MIT NER Benchmarks.
III-D Ablation
To better understand the impact of negative sampling strategies and entity type dropout on model performance, we conducted an ablation study. We analyzed varying the proportion of negative samples, the method for negative sampling, and the inclusion of random entity type dropout during training. For efficiency and consistency, all models were trained for 25K steps using the augmented NuNER dataset [9] and evaluated on the MIT NER benchmarks.
@@ III-D Ablation
Table IV shows that using full entity sets from negative examples, “random sample”, consistently outperforms random negative sampling of NER types, likely due to the contextual consistency of the selected examples. We also tested a “hard negative” approach, where negative examples are chosen based on semantic similarity to positive examples. While this method performed comparably, it introduced additional overhead in querying the dataset, hence we opt for random sampling from the training set for efficiency.
Figure 3 demonstrates that using two negative samples (approximately $66\%$ negative entity types) provides the best F1 and WER results. Figure 2 highlights the effectiveness of random entity type dropout, which helps reducing entity tags hallucinations during inference. Based on these findings, we train WhisperNER with entity type dropout and negative sampling based on the entity set from two negative examples.
IV Conclusion
@@
In this paper, we propose WhisperNER, a novel approach for joint open type named entity recognition and automatic speech recognition. Our model is designed to generalize to novel entities at inference, effectively handling a diverse range of entity types. We demonstrate that WhisperNER not only outperforms traditional pipeline-based baselines in terms of entity recognition accuracy but also achieves this with minimal to no increase in transcription errors. This work marks a step forward in the integration of speech and language processing, significantly improving both the effectiveness and precision of speech-driven applications.

### v2

@@ III-A Zero-shot Open Type NER
We first evaluate WhisperNER on the challenging open-type NER setup, in which the model is prompted at inference time with novel entity types, not observed during training. We use the three open-type NER benchmarks, namely VoxPopuli-NER, LibriSpeech-NER, and Fleurs-NER. Similar to [12], each test example is annotated with positive NER tags, while negative entities are generated by randomly sampling NER types from the full dataset. We ensure that the number of negative NER labels matches the number of positive labels, resulting in a balanced dataset. The results are presented in Table I. both WhisperNER variants outperform all baseline methods in terms of NER F1 score, with only a slight decrease in WER, while being more parameter-efficient. In comparison, both GLiNER-L and NuNER add 459M parameters each ($\sim$30% increase), while GNER adds 248M parameters ($\sim$16% increase), all on top of the Whisper large backbone.
III-B Controlling the Precision-Recall Tradeoff
Controlling the precision-recall tradeoff is essential in some domains. For example, higher recall is crucial in medical diagnostics to avoid missing critical entities. As described in Section II-B, biasing the start entity token in the logit processor enables control over the precision-recall tradeoff in entity extraction. A negative bias enhances precision but lowers recall by suppressing entities, while a strong positive bias increases recall at the possible cost of false positives. In Figure 2, we visualize our entity biasing approach using the VoxPopuli-NER dataset. As can be seen, this simple inference time modification allows significant control over the precision-recall tradeoff curve.
@@ III-B Controlling the Precision-Recall Tradeoff
III-C Ablation
@@
To better understand the impact of negative sampling strategies and entity type dropout on model performance, we conducted an ablation study. We analyzed varying the proportion of negative samples, the method for negative sampling, and the inclusion of random entity type dropout during training. For efficiency and consistency, all models were trained for 25K steps using the augmented NuNER dataset [12] and evaluated on the MIT NER benchmarks. Table II shows that using full entity sets from negative examples, “random sample”, consistently outperforms random negative sampling of NER types, likely due to the contextual consistency of the selected examples. We also tested a “hard negative” approach, where negative examples are chosen based on semantic similarity to positive examples. While this method performed comparably, it introduced additional overhead in querying the dataset, hence we opt for random sampling from the training set for efficiency. Figure 4 demonstrates that using two negative samples (approximately $66\%$ negative entity types) provides the best F1 and WER results. Figure 3 highlights the effectiveness of random entity type dropout, which helps reduce entity tag hallucinations during inference. Based on these findings, we train WhisperNER with entity type dropout and negative sampling based on the entity set from two negative examples.
Figure 5: Sequence Length Analysis: The BIO tagging scheme significantly increases token count, while the WhisperNER format keeps predicted sequence length much closer to the Whisper-only baseline.
III-D Supervised Finetuning
While the main focus of our paper is on the challenging open-type NER benchmark, some scenarios requires operating on a fixed set of entities. Here, we evaluate the supervised fine-tuning performance of our WhisperNER model using commonly used closed set NER dataset from the MIT NER benchmark [30], namely MIT-Movie and MIT-Restaurant. For baseline methods, we independently finetune both the Whisper model and the NER model. The results are summarized in Table III. Our method outperforms all baselines, achieving the best performance in both WER and NER F1 scores. These results highlight WhisperNER’s effectiveness in handling transcription and entity recognition in a supervised setting.
III-E Runtime Analysis
We analyze the predicted sequence length as a proxy for runtime and computational efficiency. Specifically, we compare WhisperNER (start/end marker format), WhisperNER-BIO (BIO tagging scheme), and the original Whisper-only baseline. Our findings presented in Figure 5 show that the BIO tagging scheme introduces a significantly higher number of tokens, increasing the overall output length and, consequently, the decoding cost. In contrast, the start/end marker format used in Whisper-NER adds substantially fewer special tokens, resulting in predicted sequences that are much closer in length to the Whisper-only baseline. This suggests that Whisper-NER is more efficient at inference time than its BIO variant, enabling cost-effective joint transcription and entity recognition.
III-F Zero-shot Language Generalization
To assess the cross-lingual capabilities of WhisperNER, we conduct a preliminary zero-shot evaluation on languages not seen during training. We apply the English-trained model to open-type NER tasks in Spanish and French using the VoxPopuli dataset. As shown in Table IV, the model exhibits some generalization ability, achieving F1 scores of 29.3 and 28.6 on Spanish and French, respectively, with relatively low word error rates. These results suggest that our approach can transfer to other languages to some extent, even without exposure during training. This points to the potential of extending WhisperNER to new languages through multilingual pretraining or lightweight fine-tuning on small labeled datasets.
TABLE IV: Zero-shot generalization to unseen languages using the English-trained WhisperNER model. Evaluated on VoxPopuli.
Language F1 $\uparrow$ WER $\downarrow$
Spanish 29.3 4.8
French 28.6 13.9
IV Limitations
While WhisperNER demonstrates strong performance in end-to-end transcription and NER, several limitations remain. First, as described in Section II, we adopt an LLM-based annotation approach to generate NER labels for ASR datasets that lack human annotations, following the method proposed in [12]. While this enables scalable and cost-effective supervision, relying on an LLM for both training and test annotations may introduce systematic biases and reduce the independence of the evaluation set. As a result, model performance may reflect alignment with the LLM’s labeling behavior rather than genuine generalization from the speech model. Second, we observe a modest degradation in word error rate, which we hypothesize is due to fine-tuning on synthetic speech generated via text-to-speech. Although TTS supports scalable data augmentation, it may introduce mismatches in acoustic characteristics when compared to natural speech. Third, our model introduces additional generation time due to the inclusion of NER tag tokens in the output sequence, which increases decoding length. While this overhead is inherent to the joint modeling strategy, future work may explore more efficient decoding or selective tagging approaches to reduce computational cost.
V Conclusion
@@
In this paper, we propose WhisperNER, a novel approach for joint open type named entity recognition and automatic speech recognition. Our model is designed to generalize to novel entities at inference, effectively handling a diverse range of entity types. We demonstrate that WhisperNER not only outperforms traditional pipeline-based baselines in terms of entity recognition accuracy but also achieves this with minimal to no increase in transcription errors. Our work marks a step forward in the integration of speech and NLP, significantly improving both the effectiveness and precision of speech applications.